why a thing is weightless inside a artificial satelite while it has weight on moon which one is also a natural satelite

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear abhishek

artificial setelite revolves around the earth near to the earth . with a acceleration of GM/R2    

 Where R is radius of orbit if setelite rotation.

when mass is inside the setellite .two force act on it gravitational pull by earth and normal reaction

so

GMm/R2 -N =m(GM/R2)  according to newton law

so N=0

 so body is weightless in this case

 

but when a body is on the moon then here gravitational pull by moon on body can not be neglected as its mass of moon is large

so here wieght of the mass due to gravitational pull by moon

F= GMmm/Rm2


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the

answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin